OpenStudy (anonymous):

Can someone please explain how to solve this? Simplify completely quantity 2 x squared plus 13 x plus 20 all over x squared minus 5 x minus 36. I know you're supposed to find the greatest common factor or find the numbers that come out to 20 and add up to 13, but what about when you cant do either?

OpenStudy (anonymous):

\[\frac{2x^2+12x+20}{x^2-5x-36}\]?

OpenStudy (anonymous):

The 12 should be a 13 :)

OpenStudy (anonymous):

\[\frac{2x^2+13 x+20}{x^2-5x-36}\]

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

idea is to factor the numerator and denominator, then cancel any common factors

OpenStudy (anonymous):

for example, the denominator is \[x^2-5x-36=(x+4)(x-9)\]

OpenStudy (anonymous):

when you factor the numerator, one of the factors will also be \(x+4\) try it then you can cancel

OpenStudy (anonymous):

Thats kinda what I'm confused on the most, factoring when theres 3 numbers in the numerator

OpenStudy (anonymous):

gonna factor like this \[2x^2+13x+20=(x+4)(2x+something)\]

OpenStudy (anonymous):

if you want a big hint, the product of the "something" and 4 has to be 20

OpenStudy (anonymous):

So 5?

OpenStudy (anonymous):

yeah 5 you can check by multiplication that \[(x+4)(2x+5)=2x^2+13x+20\]

OpenStudy (anonymous):

so you have \[\frac{(x+4)(2x+5)}{(x+4)(x-9)}\] then cancel

OpenStudy (anonymous):

So its 2x+5 over x-9?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

Is there an easy way to remember the steps to factor?

OpenStudy (anonymous):

no just have to grind it till you find it

OpenStudy (anonymous):

you can start with \[(x+a)(2x+b)\] for the first one, because you know what \(x\times 2x=2x^2\) and you will also know that \(ab=20\) but there are several ways to factor 20 just turns out 4 and 5 work for this one

OpenStudy (anonymous):

Ah alright, thank you so much :)